You know that overall there is a 60% chance to observe birds engaging in a specific mating dance. You want to study this dance, so you need to know how many birds to observe to have a good chance to see it. How many birds would you need to observe to have a 95% chance of observing the dance?

Let’s say your previous research has also shown that the same species in Florida dances for 8 minutes on average, with a standard deviation of 1.5 minutes. How many birds would you need to observe to get a good estimate of the dancing time in South Texas, plus of minus a standard error of 30 seconds?
Answer #1

Standard Deviation ,   σ =    1.5 minutes   
sampling error ,    E = 30 seconds = 0.5 minutes
Confidence Level ,   CL=   95%                  
                          
alpha =   1-CL =   5%                  
Z value =    Zα/2 =    1.960   [excel formula =normsinv(α/2)]              
                          
Sample Size,n = (Z * σ / E )² = (   1.960   *   1.5   /   0.5   ) ² =   34.573
                          
                          
So,Sample Size needed=       35